OpenStudy (anonymous):

A tank can be filled by one pipe in four hours and by a second pipe in six hours; and when it is full, the tank can be drained by a third pipe in three hours. If the tank is empty and all three pipes are open, in how many hours will the tank be filled?

OpenStudy (anonymous):

let the tank be filled in x hours ,then \[\frac{ x }{ 4 }+\frac{ x }{ 6 }-\frac{ x }{ 3 }=1\] L.C.M. of 4,6,3=12 multiply by 12 3x+2x-4x=12 x=?